OpenStudy (jiteshmeghwal9):

Prove :- sin(180-a)=sina cos(360-a)=cosa tan(180+a)=tana

OpenStudy (mayankdevnani):

use sin(a-b) formula !

OpenStudy (mayankdevnani):

\[\large \bf \sin(a-b)=\sin{a}\cos{b}-\sin{b}\cos{a}\] Put a=180 and b=a

OpenStudy (denonakavro):

For sine check here: http://math.stackexchange.com/questions/569105/why-is-sinx-sin180-circ-x

OpenStudy (denonakavro):

For cos Since cos(a-b) = cosacosb + sinasinb Thus cos(360 - x) = cos360*cosx + sin360*sinx with cos360=1 and sin360=0, thus cos(360 - x) = cosx

OpenStudy (mayankdevnani):

understood jitesh ?

OpenStudy (jiteshmeghwal9):

yes i gt it but how can we gt the general expression for sine i.e; \[\LARGE{n \pi +(-1)^n \alpha}\]